# Multilevel lattice codes from Hurwitz quaternion integers

### v2

@@ IV-A Decoding complexity
5
$\mathscr{H}$
$(1+2i)\cdot(1-2i)$
$\mathbb{Z}[\omega]$
$(1+3\omega)\cdot(-2-3\omega)$
7
$\mathscr{H}$
$(1+i+j+2k)\cdot(1-i-j-2k)$
11
$\mathscr{H}$
$(1+i+3j)\cdot(1-i-3j)$
@@
$(2+3i)\cdot(2-3i)$
$\mathbb{Z}[\omega]$
$(1+4\omega)\cdot(-3-4\omega)$
13
$\mathscr{H}$
$(1+2i+2j+2k)\cdot(1-2i-2j-2k)$
$\mathbb{Z}[i]$
$3\cdot(1+2i)\cdot(1-2i)$
15
$\mathscr{H}$
$(1+i+j)(1-i-j)(1+2i)(1-2i)$
$\mathbb{Z}[i]$
17
$\mathscr{H}$
$(1+4i)\cdot(1-4i)$
$\mathbb{Z}[\omega]$
$(2+5\omega)\cdot(-3-5\omega)$
19
$\mathscr{H}$
$(1+i+j+4k)\cdot(1-i-j-4k)$
21
$\mathscr{H}$
$(1+i+j)(1-i-j)(1+i+j+2k)(1-i-j-2k)$
23
$\mathscr{H}$
$(1+2i+2j+3k)\cdot(1-2i-2j-3k)$
$\mathbb{Z}[i]$
$(2+5i)\cdot(2-5i)$
29
$\mathscr{H}$
$(2+3i+4j)\cdot(2-3i-4j)$
$\mathbb{Z}[\omega]$
$(1+6\omega)\cdot(-5-6\omega)$
31
$\mathscr{H}$
$(1+i+2j+5k)\cdot(1-i-2j-5k)$
33
$\mathscr{H}$
$(1+i+j)(1-i-j)(1+i+3j)(1-i-3j)$
$\mathbb{Z}[i]$
$7\cdot(1+2i)\cdot(1-2i)$
$\mathbb{Z}[\omega]$
$5\cdot(1+3\omega)\cdot(-2-3\omega)$
35
$\mathscr{H}$
$(1+2i)(1-2i)(1+i+j+2k)(1-i-j-2k)$
$\mathbb{Z}[i]$
$(1+6i)\cdot(1-6i)$
$\mathbb{Z}[\omega]$
$(3+7\omega)\cdot(-4-7\omega)$
37
$\mathscr{H}$
$(1+2i+4j+4k)\cdot(1-2i-4j-4k)$
$\mathbb{Z}[i]$
$3\cdot(2+3i)\cdot(2-3i)$
39
$\mathscr{H}$
$(1+i+j)(1-i-j)(1+2i+2j+2k)(1-2i-2j-2k)$
Figure 1: Comparison of decoding complexity for constellations with cardinality $q^{4}$ in $\mathbb{R}^{4}$ obtained through Construction $\pi_{A}$ over $\mathbb{Z}[i]$, $\mathbb{Z}[\omega]$, and $\mathscr{H}$. Here, $|\mathscr{C}_{\text{max}}|$ denotes the maximum cardinality among the layers in the encoding.
Constellations of the Hurwitz integers have been considered in the context of four dimension coherent optical systems [18, 19, 45]. It has been shown that they can have good performance regarding shaping gain and different constrains for the shape such the hypercube, Voronoi region and spherical are analysed [21, 46, 47, 26]. Particularly constellations of $2^{b}$ points are studied. It seems to be interesting to investigate if the constellations $(\mathscr{H}/q\mathscr{H})^{n}$ of $q^{4n}$ points, $q=p_{1}\cdot\dots\cdot p_{r},\ p_{i}$ prime, considered above, may enlarge this scenario also providing good shaping gain with lower computational complexity.
@@
$\left(\frac{\mathscr{H}}{\mathscr{H}(1+2i)}\right)^{2}\times\left(\frac{\mathscr{H}}{\mathscr{H}(1-2i)}\right)^{2}\rightarrow\left(\frac{\mathscr{H}}{5\mathscr{H}}\right)^{2}$
and
$\left(\frac{\mathbb{Z}[i]}{\mathbb{Z}[i](1+2i)}\right)^{4}\times\left(\frac{\mathbb{Z}[i]}{\mathbb{Z}[i](1-2i)}\right)^{4}\rightarrow\left(\frac{\mathbb{Z}[i]}{5\mathbb{Z}[i]}\right)^{4}$
[…]
@@ V Goodness for channel coding
The main objective of this section is to demonstrate that Construction $\pi_{A}$ lattice from codes over $\mathbb{Z},\mathbb{Z}[i],\mathbb{Z}[\omega]$ and $\mathscr{H}$ are good for AWGN channel coding. In [1] a different proof for the three first cases is presented.
@@ V-A Balanced set of codes over rings
Let $\mathscr{R}$ be a finite ring, and $\mathscr{R}^{\times}$ be its units. Denote by $(\mathscr{R}^{n})^{*}$ the set of vectors in $\mathscr{R}^{n}$ such that at least one coordinate is a unit.
In this section, we consider a linear code $\mathscr{C}\subset\mathscr{R}^{n}$ as a free [^1] The usual definition of linear codes over rings involves considering $\mathscr{C}$ as a $\mathscr{R}-$submodule. However, in this section, we require linear codes with a basis, and consequently, we need to restrict the codes to the case of free submodules. These codes are also called free linear codes [48]. $\mathscr{R}$-submodule of $\mathscr{R}^{n}$. The rank $k$ of such linear code $\mathscr{C}$, $\text{rank}(\mathscr{C})=k$, is the number of elements of a basis of $\mathscr{C}$ [49]. We define a balanced set of codes as follows.
@@
Consider a non-empty set of linear codes $\mathscr{C}_{b}$ of the same cardinality. We say that $\mathscr{C}_{b}$ is balanced in $\mathscr{R}^{n}$ if every $\mathbf{x}\in(\mathscr{R}^{n})^{*}$ belongs to the same number, $L$, of codes of $\mathscr{C}_{b}$.
As an example, let $\mathscr{R}=\mathbb{Z}_{p}$, $p$ prime. Since all non vanishing element of $\mathbb{Z}_{p}$ is a unit, $(\mathbb{Z}_{p}^{n})^{*}=\mathbb{Z}_{p}^{n}-\{\mathbf{0}\}$. Therefore,
$|(\mathbb{Z}_{p}^{n})^{*}|=p^{n}-1,$
Clearly, the set $\mathscr{C}_{k}$ of all codes with rank $k$ over $\mathbb{Z}_{p}^{n}$ is a balanced set with
$L\cdot(p^{n}-1)=|\mathscr{C}_{k}|\cdot(p^{k}-1),$
where $|\mathscr{C}_{k}|$ is the number of codes in $\mathscr{C}_{k}$, [14]. In fact, $|\mathscr{C}_{k}|$ is the so called Gaussian binomial coefficient,
$|\mathscr{C}_{k}|=\binom{n}{k}_{p}=\dfrac{(p^{n}-1)(p^{n}-p)\dots(p^{n}-p^{k-1})}{(p^{k}-1)(p^{k}-p)\dots(p^{k}-p^{k-1})}.$
For a general finite ring $\mathscr{R}$ with $q$ elements, we have that the set $\mathscr{C}_{k}$ of all linear codes with rank $k$ is balanced with $|\mathscr{C}|=M$ for $\mathscr{C}\in\mathscr{C}_{k}$ and
$L\cdot|(\mathscr{R}^{n})^{*}|\leq|\mathscr{C}_{k}|\cdot(M-1),$ (11)
because the number of elements of $(\mathscr{R}^{n})^{*}$ in each $\mathscr{C}\in\mathscr{C}_{k}$ is at most $M-1\leq q^{k}-1$, [50].
@@ V-B Balanced set of codes from the Chinese remainder theorem
$\varphi:(R/\cap_{j=1}^{s}I_{j})^{n}\rightarrow(R/I_{1})^{n}\times...\times(R/I_{s})^{n}.$ (12)
@@
where, $\mathscr{C}\subseteq(R/\cap_{j=1}^{s}I_{j})^{n}$, and $\mathscr{C}_{j}\subseteq(R/I_{j})^{n},j=1,\ldots,s$.
@@
Let $\mathscr{C}_{k_{1}},\mathscr{C}_{k_{2}},\ldots,\mathscr{C}_{k_{s}}$ be the balanced sets of all codes of rank $k_{j}$ over the finite rings $(R/I_{j})^{n},$ where $R=\mathbb{Z},\mathbb{Z}[i],\mathbb{Z}[\omega]$ or $\mathscr{H}$ and $I_{j},j=1,\ldots,s$ are relatively prime ideals in $R$. Then, it follows that
$\mathscr{C}_{k}=\varphi^{-1}(\mathscr{C}_{k_{1}}\times\mathscr{C}_{k_{2}}\times\ldots\times\mathscr{C}_{k_{s}}),$ (13)
@@
Proof: A codes $\mathscr{C}$ in $\mathscr{C}_{k}$ has a cardinality $|\mathscr{C}|=|\mathscr{C}_{1}|\cdot|\mathscr{C}_{2}|\cdot\ldots\cdot|\mathscr{C}_{s}|=q_{1}^{k_{1}}\cdot q_{2}^{k_{2}}\cdot\ldots\cdot q_{s}^{k_{s}}$, where $\mathscr{C}_{j}\in\mathscr{C}_{k_{j}},j=1,...,s$ and $k=\text{rank}(\mathscr{C})=\max\{\text{rank}(\mathscr{C}_{j})\}$, [48].
Let $\mathscr{R}^{n}=(R/\cap_{j=1}^{s}I_{j})^{n}$. For any element $\mathbf{x}$ in $(\mathscr{R}^{n})^{*}$, we can find a linear bijection $T:\mathscr{R}^{n}\rightarrow\mathscr{R}^{n}$ such that $T(\mathbf{x})=(1,0,\ldots,0)=\mathbf{e}_{1}$. Indeed, taking $\mathbf{x}=(x_{1},\ldots,x_{j},\ldots,x_{n})\in(\mathscr{R}^{n})^{*}$ with the $j-$th coordinate as a unit. A basis $\{\mathbf{x},\mathbf{w}_{1},...,\mathbf{w}_{n-1}\}$ including $\mathbf{x}$ can be chosen, allowing us to define a linear bijection $T(\mathbf{x})=\mathbf{e}_{1}$.
Since $T$ preserves rank and $\mathbf{x}\in\mathscr{C}$ if and only if $\mathbf{e}_{1}\in T(\mathscr{C})$ then $\mathbf{x}$ belongs to the same number of codes in $\mathscr{C}_{k}$ as $\mathbf{e}_{1}\in T(\mathscr{C})$, where $\mathscr{C}\in\mathscr{C}_{k}$. Now, we can express $\mathbf{e}_{1}=\varphi^{-1}(\mathbf{e}_{1}^{(1)},\mathbf{e}_{1}^{(2)},\ldots,\mathbf{e}_{1}^{(s)})$, where $\mathbf{e}_{1}^{(j)}=(1,0,...,0)$ belongs to $L_{j}$ codes of $\mathscr{C}_{k_{j}}$ for $j=1,\ldots,s$. Thus, $\mathbf{e}_{1}$ belongs to $L=L_{1}\cdot L_{2}\cdot\ldots\cdot L_{s}$ codes of $\mathscr{C}_{k}$, confirming that (13) indeed defines a balanced set of codes in $\mathscr{R}^{n}$.
@@
We will refer to the set of codes (13) as a balanced set of Construction $\pi_{A}$ codes.
@@
Let $\mathscr{R}$ be a finite ring (not necessarily commutative), $g:\mathscr{R}^{n}\rightarrow\mathbb{R}^{+}$ be a function and $\mathscr{C}_{k}$ be a balanced set of codes. Then,
$\mathbb{E}_{\mathscr{C}_{k}}\left[\sum_{c\in\mathscr{C}\cap(\mathscr{R}^{n})^{*}}g(c)\right]\leq\dfrac{|\mathscr{R}|^{k}-1}{|(\mathscr{R}^{n})^{*}|}\sum_{\mathbf{x}\in(\mathscr{R}^{n})^{*}}g(\mathbf{x}),$ (14)
where the expectation is with respect to the uniform distribution on $\mathscr{C}_{k}$.

### v3

@@ 3 $\mathscr{H}$ $(1+i+j)\cdot(1-i-j)$ $3^{2}$
5 $\mathscr{H}$ $(1+2i)\cdot(1-2i)$ $5^{2}$
$\mathbb{Z}[\omega]$ $(1+3\omega)\cdot(-2-3\omega)$
7 $\mathscr{H}$ $(1+i+j+2k)\cdot(1-i-j-2k)$ $7^{2}$
11 $\mathscr{H}$ $(1+i+3j)\cdot(1-i-3j)$ $11^{2}$
$\mathbb{Z}[i]$ $(2+3i)\cdot(2-3i)$
$\mathbb{Z}[\omega]$ $(1+4\omega)\cdot(-3-4\omega)$
13 $\mathscr{H}$ $(1+2i+2j+2k)\cdot(1-2i-2j-2k)$ $13^{2}$
$\mathbb{Z}[i]$ $3\cdot(1+2i)\cdot(1-2i)$ $3^{4}$
15 $\mathscr{H}$ $(1+i+j)(1-i-j)(1+2i)(1-2i)$ $5^{2}$
@@
17 $\mathscr{H}$ $(1+4i)\cdot(1-4i)$ $17^{2}$
$\mathbb{Z}[\omega]$ $(2+5\omega)\cdot(-3-5\omega)$
19 $\mathscr{H}$ $(1+i+j+4k)\cdot(1-i-j-4k)$ $19^{2}$
$(1+i+j)(1-i-j)(1+i+j+2k)$
21 $\mathscr{H}$ $(1-i-j-2k)$ $21^{2}$
23 $\mathscr{H}$ $(1+2i+2j+3k)\cdot(1-2i-2j-3k)$ $23^{2}$
$\mathbb{Z}[i]$ $(2+5i)\cdot(2-5i)$
29 $\mathscr{H}$ $(2+3i+4j)\cdot(2-3i-4j)$ $29^{2}$
$\mathbb{Z}[\omega]$ $(1+6\omega)\cdot(-5-6\omega)$
31 $\mathscr{H}$ $(1+i+2j+5k)\cdot(1-i-2j-5k)$ $31^{2}$
$(1+i+j)(1-i-j)(1+i+3j)$
33 $\mathscr{H}$ $(1-i-3j)$ $11^{2}$
$\mathbb{Z}[i]$ $7\cdot(1+2i)\cdot(1-2i)$ $7^{4}$
$\mathbb{Z}[\omega]$ $5\cdot(1+3\omega)\cdot(-2-3\omega)$ $5^{4}$
$(1+2i)(1-2i)(1+i+j+2k)$
35 $\mathscr{H}$ $(1-i-j-2k)$ $7^{2}$
$\mathbb{Z}[i]$ $(1+6i)\cdot(1-6i)$
$\mathbb{Z}[\omega]$ $(3+7\omega)\cdot(-4-7\omega)$
37 $\mathscr{H}$ $(1+2i+4j+4k)\cdot(1-2i-4j-4k)$ $37^{2}$
$\mathbb{Z}[i]$ $3\cdot(2+3i)\cdot(2-3i)$
$(1+i+j)(1-i-j)(1+2i+2j+2k)$
39 $\mathscr{H}$ $(1-2i-2j-2k)$ $13^{2}$
@@
$\left(\frac{\mathscr{H}}{\mathscr{H}(1+2i)}\right)^{2}\times\left(\frac{\mathscr{H}}{\mathscr{H}(1-2i)}\right)^{2}\rightarrow\left(\frac{\mathscr{H}}{5\mathscr{H}}\right)^{2},$
$\left(\frac{\mathbb{Z}[i]}{\mathbb{Z}[i](1+2i)}\right)^{4}\times\left(\frac{\mathbb{Z}[i]}{\mathbb{Z}[i](1-2i)}\right)^{4}\rightarrow\left(\frac{\mathbb{Z}[i]}{5\mathbb{Z}[i]}\right)^{4},$
and codes $\mathscr{C}_{1}\in(\mathscr{H}/\mathscr{H}(1+2i))^{2}$ and $\mathscr{C}_{2}\in(\mathscr{H}/\mathscr{H}(1-2i))^{2}$ with $\text{rank}(\mathscr{C}_{1})=1$ and $\text{rank}(\mathscr{C}_{2})=2$, respectively, then using Construction $\pi_{A}$, we have a code $\mathscr{C}\subset(\mathscr{H}/5\mathscr{H})^{2}$ with cardinality $|\mathscr{C}|=5^{6}$, and the decoding complexity is of order $5^{4}\log_{2}5^{4}$. Conversely, a code $\mathscr{C}\subset(\mathbb{Z}[i]/5\mathbb{Z}[i])^{4}$ with $|\mathscr{C}|=5^{6}$ can be obtained by Construction $\pi_{A}$ over the Gaussian integers from codes with rank $3$ in each layer and therefore the decoding complexity is of order $5^{3}\log_{2}5^{3}$.
Constellations of the Hurwitz integers have been considered in the context of four-dimensional coherent optical systems [17, 18, 45]. It has been shown that they can have good performance regarding shaping gain and different constrains for the shape such the hypercube, Voronoi region and spherical [20, 46, 47, 25]. Particularly constellations of $2^{b},b\in\mathbb{N}$ points are studied.
To evaluate the performance of existing constellations, Table III compares some parameters: the minimum distance $d_{\min}$, the average symbol energy ($E$), the normalized second moment (NSM), and the constellation figure of merit (CFM). In this comparison, it is considered Erik Agrell’s database of sphere packings and constellations [48]. For clarity, we briefly define each parameter below.
The average symbol energy $E$ quantifies the mean energy required to transmit a symbol and is defined as $E=\frac{1}{M}\sum_{i=1}^{M}\|x_{i}\|^{2},$ where $M$ denotes the number of points in the constellation, and $x_{i}\in\mathbb{R}^{n}$ is the vector representing the $i-$th constellation point. This quantity corresponds to the mean squared Euclidean norm of the constellation points.
The normalized second moment (NSM) measures the energy efficiency of a constellation relative to its minimum distance and it is given by, $\text{NSM}=\frac{E}{d_{\min}^{2}},$ where $d_{\min}$ is the minimum Euclidean distance between any two distinct constellation points. Lower NSM values indicate tighter packings.
The constellation figure of merit (CFM) [49] serves as a quality metric that balances minimum distance and average energy. It is defined as $\text{CFM}=\frac{2d_{\min}^{2}}{E}.$ and reflects how well-spread a constellation is with respect to energy usage. For an easy comparison, we report CFM in decibels (dB), computed as $10\log_{10}\text{CFM}$.
Constellations and sphere packings can be computed under many criteria depending on the communication scenario. When comparing constellations of the same dimension and cardinality, the one with lowest NSM is preferred, as it indicates higher energy efficiency. Alternatively, when comparing modulation formats under the same bandwidth, the format with the largest CFM offers better performance in terms of distance-to energy balance.
In Table III, we adopt the notation used in [48] to refer to various classes of sphere packings in dimension 4. The entries w4_* correspond to good packings of selected sizes originally proposed by Welti in [20], obtained by selecting points from $D_{4}$ lattice that lie in certain chosen hyperspheres. The labels l4_* refer to optimal subsets of the $D_{4}$ lattice (in terms of constellation design), constructed in unpublished work by Agrell. These subsets are not necessarily optimal sphere packings. Finally, the c4_* entries represent conjectured optimal sphere packings drawn from G. Nebe and N. Sloane unpublished tables [50].
TABLE III: Comparison of Average Symbol Energy, NSM, and CFM for Construction $\pi_{A}$ lattice codes via Hurwitz integers and existing approaches with low cardinality.
Code size $d_{\min}$ $E$ NSM CFM (dB)
tetra4 $1.732$ $2.666$ $0.889$ $3.521$
l4_9 $18$ $192$ $0.592$ $5.282$
c4_9 $1.608$ $1.481$ $\mathbf{0.572}$ $\mathbf{5.433}$
$\mathscr{H}_{\pi_{A}}$ $9$ $1.732$ $2.666$ $0.889$ $3.521$
2prism4 $1.175$ $2$ $1.447$ $1.404$
l4_25 $2$ $3.84$ $\mathbf{0.96}$ $\mathbf{3.187}$
c4_25 $2$ $3.84$ $\mathbf{0.96}$ $\mathbf{3.187}$
$\mathscr{H}_{\pi_{A}}$ $25$ $2.236$ $4.8$ $\mathbf{0.96}$ $\mathbf{3.187}$
2prism4 $0.867$ $2$ $2.655$ $-1.231$
w4_49 $1.414$ $2.938$ $\mathbf{1.469}$ $\mathbf{1.338}$
l4_49 $2$ $5.877$ $\mathbf{1.469}$ $\mathbf{1.338}$
$\mathscr{H}_{\pi_{A}}$ $49$ $2.645$ $10.285$ $\mathbf{1.469}$ $\mathbf{1.338}$
l4_81 $162$ $49712$ $\mathbf{1.894}$ $\mathbf{0.235}$
$\mathscr{H}_{\pi_{A}}$ $81$ $1$ $2.074$ $2.074$ $-0.157$
2prism4 $0.563$ $2$ $6.299$ $-4.982$
l4_121 $242$ $135936$ $\mathbf{2.321}$ $\mathbf{-0.646}$
$\mathscr{H}_{\pi_{A}}$ $121$ $3.316$ $26.181$ $2.380$ $-0.755$
w4_169 $1$ $5.396$ $\mathbf{2.698}$ $\mathbf{-1.300}$
$\mathscr{H}_{\pi_{A}}$ $169$ $3.605$ $36.92$ $2.840$ $-1.523$
@@ V Goodness for channel coding
The main objective of this section is to show that Construction $\pi_{A}$ lattices from codes over $\mathbb{Z},\mathbb{Z}[i],\mathbb{Z}[\omega]$ and $\mathscr{H}$ can be “good for coding” over the AWGN channel. In [3], a proof with different approach, using arguments more closely to those of Forney [28], is presented for $\mathbb{Z},\mathbb{Z}[i]$ and $\mathbb{Z}[\omega]$.
@@ V-A Balanced set of codes over rings
Let $\mathscr{R}$ be a finite ring, and $\mathscr{R}^{\times}$ be its units. Denote by $(\mathscr{R}^{n})^{\times}$ the set of vectors in $\mathscr{R}^{n}$ such that at least one coordinate is a unit.
In this section, we consider a linear code $\mathscr{C}\subset\mathscr{R}^{n}$ as a free [^2] In the previous sections we have assumed the usual definition of linear codes over rings as a $\mathscr{R}-$submodule. However, in this section, we require also that a linear code has a basis. These codes are called free linear codes [51]. $\mathscr{R}$-submodule of $\mathscr{R}^{n}$. The rank $k$ of such a linear code $\mathscr{C}$, $\text{rank}(\mathscr{C})=k$, is the number of elements of a basis of $\mathscr{C}$ [52]. We define a balanced set of codes as follows.
@@
Consider a non-empty set of linear codes $\mathscr{C}_{b}$ of the same cardinality. We say that $\mathscr{C}_{b}$ is balanced in $\mathscr{R}^{n}$ if every $x\in(\mathscr{R}^{n})^{\times}$ belongs to the same number, $L$, of codes of $\mathscr{C}_{b}$.
As an example, let $\mathscr{R}=\mathbb{Z}_{p}$, $p$ prime. Since a nonvanishing element of $\mathbb{Z}_{p}$ is a unit, $(\mathbb{Z}_{p}^{n})^{\times}=\mathbb{Z}_{p}^{n}-\{\mathbf{0}\}$. Therefore,
$|(\mathbb{Z}_{p}^{n})^{\times}|=p^{n}-1,$
Clearly, the set $\mathscr{C}_{b_{k}}$ of all linear codes with rank $k$ over $\mathbb{Z}_{p}^{n}$ is a balanced set with
$L\cdot(p^{n}-1)=|\mathscr{C}_{b_{k}}|\cdot(p^{k}-1),$
where $|\mathscr{C}_{b_{k}}|$ is the number of codes in $\mathscr{C}_{b_{k}}$, [27]. In fact, $|\mathscr{C}_{b_{k}}|$ is the so called Gaussian binomial coefficient
$|\mathscr{C}_{b_{k}}|=\binom{n}{k}_{p}=\dfrac{(p^{n}-1)(p^{n}-p)\dots(p^{n}-p^{k-1})}{(p^{k}-1)(p^{k}-p)\dots(p^{k}-p^{k-1})}.$
For a general finite ring $\mathscr{R}$ with $q$ elements, we have that the set $\mathscr{C}_{b_{k}}$ of all linear codes with rank $k$ is balanced with $|\mathscr{C}|=M$ for $\mathscr{C}\in\mathscr{C}_{b_{k}}$ and
$L\cdot|(\mathscr{R}^{n})^{\times}|\leq|\mathscr{C}_{b_{k}}|\cdot(M-1)$ (11)
because the number of elements of $(\mathscr{R}^{n})^{\times}$ in each $\mathscr{C}\in\mathscr{C}_{k}$ is at most $M-1\leq q^{k}-1$, [53].
@@ V-B Balanced set of codes from the Chinese Remainder Theorem
$\varphi:(R/\cap_{j=1}^{s}I_{j})^{n}\rightarrow(R/I_{1})^{n}\times\ldots\times(R/I_{s})^{n}.$ (12)
@@
where, $\mathscr{C}\subset(R/\cap_{j=1}^{s}I_{j})^{n}$, and $\mathscr{C}_{j}\subset(R/I_{j})^{n},j=1,\ldots,s$.
@@
Let $\mathscr{C}_{b_{k_{1}}},\mathscr{C}_{b_{k_{2}}},\ldots,\mathscr{C}_{b_{k_{s}}}$ be the balanced sets of all codes of rank $k_{j}$ over the finite rings $(R/I_{j})^{n},$ where $R=\mathbb{Z},\mathbb{Z}[i],\mathbb{Z}[\omega]$ or $\mathscr{H}$ and $I_{j},j=1,\ldots,s$ are relatively prime ideals in $R$. Then, it follows that
$\mathscr{C}_{b_{k}}=\varphi^{-1}(\mathscr{C}_{b_{k_{1}}}\times\mathscr{C}_{b_{k_{2}}}\times\ldots\times\mathscr{C}_{b_{k_{s}}}),$ (13)
@@
Proof: A code $\mathscr{C}$ in $\mathscr{C}_{b_{k}}$ has cardinality $|\mathscr{C}|=|\mathscr{C}_{1}|\cdot|\mathscr{C}_{2}|\cdot\ldots\cdot|\mathscr{C}_{s}|=q_{1}^{k_{1}}\cdot q_{2}^{k_{2}}\cdot\ldots\cdot q_{s}^{k_{s}}$, where $\mathscr{C}_{j}\in\mathscr{C}_{b_{k_{j}}},j=1,\ldots,s$ and $k=\text{rank}(\mathscr{C})=\max\{\text{rank}(\mathscr{C}_{j})\}$, [51].
Let $\mathscr{R}^{n}=(R/\cap_{j=1}^{s}I_{j})^{n}$. For any element $x$ in $(\mathscr{R}^{n})^{\times}$, we can find a linear bijection $T:\mathscr{R}^{n}\rightarrow\mathscr{R}^{n}$ such that $T(x)=(1,0,\ldots,0)=\mathbf{e}_{1}$. Indeed, taking $x=(x_{1},\ldots,x_{j},\ldots,x_{n})\in(\mathscr{R}^{n})^{\times}$ with the $j-$th coordinate as a unit, a basis $\{x,z_{1},\ldots,z_{n-1}\}$ including $x$ can be chosen, allowing us to define a linear bijection with $T(x)=\mathbf{e}_{1}$.
Since $T$ preserves rank and $x\in\mathscr{C}$ if and only if $\mathbf{e}_{1}\in T(\mathscr{C})$ then $x$ belongs to the same number of codes in $\mathscr{C}_{b_{k}}$ which is the number of codes $T(\mathscr{C})\in\mathscr{C}_{b_{k}}$ that contains $\mathbf{e}_{1}$. Now, we can express $\mathbf{e}_{1}=\varphi^{-1}(\mathbf{e}_{1}^{(1)},\mathbf{e}_{1}^{(2)},\ldots,\mathbf{e}_{1}^{(s)})$, where $\mathbf{e}_{1}^{(j)}=(1,0,\ldots,0)$ belongs to $L_{j}$ codes of $\mathscr{C}_{b_{k_{j}}}$ for $j=1,\ldots,s$. Thus, $\mathbf{e}_{1}$ belongs to $L=L_{1}\cdot L_{2}\cdot\ldots\cdot L_{s}$ codes of $\mathscr{C}_{b_{k}}$, confirming that (13) indeed defines a balanced set of codes in $\mathscr{R}^{n}$.
@@
We will refer to the above set of codes $\mathscr{C}_{b_{k}}$ as the balanced set of Construction $\pi_{A}$ codes from (13).
@@
Let $\mathscr{R}$ be a finite ring (not necessarily commutative), $g:\mathscr{R}^{n}\rightarrow\mathbb{R}_{+}$ be a function and $\mathscr{C}_{b_{k}}$ be a balanced set of $\mathscr{R}-$linear codes of rank $k$. Then,
$\mathbb{E}_{\mathscr{C}_{b_{k}}}\left[\sum_{c\in\mathscr{C}\cap(\mathscr{R}^{n})^{\times}}g(c)\right]\leq\dfrac{|\mathscr{R}|^{k}-1}{|(\mathscr{R}^{n})^{\times}|}\sum_{x\in(\mathscr{R}^{n})^{\times}}g(x),$ (14)
where $\mathscr{C}\in\mathscr{C}_{b_{k}}$ and the expectation is with respect to the uniform distribution on $\mathscr{C}_{b_{k}}$.